OpenStudy (anonymous):

Explain how you would find the exact value of 3/√8

OpenStudy (unklerhaukus):

use 8=2^3=2\(\times\)2^2 and √(mn)=√m√n

OpenStudy (unklerhaukus):

after that multiply by √2/√2